assignment is due at 11:45 PM on 10/07/2018. Any work submited between that time and 11:45 PM on 10/08/2018 will result in an a utomatic 10 Antonio, please try the problem again. Try 1: X 4.58s Question 9 A buffer solution is prepared by mixing 45.0 mL of 1.20 M Acetic Acid and 30.0 ml of 1.25 M Nac2H302. Determine the pH of the solution after the addition of 0.01 moles NaOH (assume there is no change in volume when the NaOH is added) Kb C2H302 5.556E-10 Your answer, 4.585, i is incorrect by less than 10%-verify that the numbers entered into your calculator are the same as those listed 4,585 Submit Problem Quick Links: 112121,ΙΣΙ6izi819110111 I 12112114 115116 1 Cancel to Assigoment Menu Back to Main Menu
Answer #1

9)

millimoles of acetic acid = 45 x 1.20 = 54

millimoles of sodium acetate = 30 x 1.25 = 37.5

millimoles of NaOH = 10 (C)

pH = pKa + log [conjugate base + C / acid - C]

    = 4.74 + log [37.5 + 10 / 54 - 10]

pH = 4.77